Consider A = <0,4,2> , B = <6,-1,0> , and C = <3,0,1>. Find scalars a, b and c such that aA + bB = (c - 1)C.


.



1
Expert's answer
2022-04-25T13:36:09-0400

"A = 
\begin{bmatrix} 
0\\
4\\
2
\end{bmatrix},
B = \begin{bmatrix} 
6\\
-1\\
0
\end{bmatrix},
C =
\begin{bmatrix} 
3\\
0\\
1
\end{bmatrix}"

"aA +bB = (c-1)C"

"a,b,c - ?"

"a\begin{bmatrix} 
0\\
4\\
2
\end{bmatrix}
+
b \begin{bmatrix} 
6\\
-1\\
0
\end{bmatrix}
= (c-1) \begin{bmatrix} 
3\\
0\\
1
\end{bmatrix}"


"\begin{cases}
6b = 3c-3 \\
4a-b = 0\\
2a = c-1
\end{cases}" "\begin{cases}
2b = c-1\\
4a = b\\
2a = c-1
\end{cases}"

From (1) and (3) "2a = 2b \rightarrow a = b"

From (2) "4a = a \rightarrow 3a = 0 \rightarrow a =0 \rightarrow b = 0"

From (3) "c-1 = 2a \rightarrow c-1 = 0 \rightarrow c =1"

Answer: "a = 0, b = 0, c = 1"